### v2

# arXiv:2602.19625v2[math.PR] 03 Mar 2026

Abstract.
Intermittent demand fluctuations pose significant challenges in disaster logistics and medical supply systems. In this study, we formulate cumulative demand as a generalized Lévy process composed of a drift term, Poisson jumps, and compound Poisson jumps, and analyze a continuous-time inventory model. The proposed framework provides a unified formulation that encompasses both drifted Poisson processes and drifted compound Poisson processes. From a mathematical perspective, we treat the reorder time as a first-passage problem of a subordinator and derive its Laplace transform via the Laplace exponent. In particular, for the drifted Poisson case, we obtain an explicit representation of the inverse Laplace exponent using the Lambert W function, which yields an analytic expression for the Laplace transform of the first-passage time. Furthermore, when the jump sizes follow exponential and Gamma distributions, we derive explicit formulas for the mean and variance of the reorder times, thereby clarifying the moment structure of first-passage times for generalized Lévy demand processes. From an operations research perspective, we explicitly characterize the expected total cost over a finite time horizon based on the distribution of cumulative demand. This study presents an analytical framework that integrates first-passage theory of Lévy processes with continuous-time inventory control.
1. Introduction
Inventory control problems under uncertain demand constitute an important research area in both probability theory and operations research. In particular, in situations such as disaster response or intermittent demand, continuous demand models are often insufficient. Lévy processes provide a well-behaved class of stochastic processes capable of modeling discontinuous changes. ( [9][6][10][5][4])
For example, Noba and Yamazaki (2025) [7] propose an inventory control policy based on Lévy processes and conduct analytical investigations. In their framework, the total cost incurred in inventory control is expressed as a general functional form, and the optimal replenishment strategy is characterized. However, difficulties remain in performing numerical computations and explicit calculations under concrete specifications.
On the other hand, Koide et al. (2024) [3] and Koide et al. (2025) [5] propose inventory control models driven by Lévy processes. By imposing engineering-oriented assumptions that facilitate computation and simulation, they explicitly derive the expected total cost and enable numerical simulations.
In this study, we consider $\{D_{t}\}_{t\geq 0}$ as a stochastic process representing cumulative demand. In particular, we focus on non-decreasing Lévy processes (subordinators), such as a drifted Poisson process (cf. the setting in Koide [5]). Let $a>0$, $Q>0$, and $n\in\mathbb{N}$. The first passage time of $D_{t}$ is defined by
$T_{n}=\inf\left\{s>0\,\middle|\,D_{s}\geq a+(n-1)Q\right\}.$
The random variable $T_{n}$ represents the time of the $n$-th replenishment. Let $x>0$ denote the initial inventory level, where xax-a is the reorder point and $Q>0$ is the order quantity. We consider a fixed-order-quantity policy. Then, the inventory level at time $t$, denoted by $X_{t}$, is given by
$X_{t}=x-D_{t}+QR_{t}\quad\text{, where}\quad R_{t}=\sum_{n\geq 1}1_{\{T_{n}<t\}}=\inf\{k\mid T_{k}>t\}.$
Here, $\{R_{t}\}_{t\geq 0}$ represents the stochastic process describing the cumulative number of orders placed up to time $t$.
In order to compute the expectation and variance of $X_{t}$, as well as the expected total cost, it is necessary to evaluate the expectation and variance of $T_{n}$. However, the distribution of $T_{n}$ is generally unknown, making direct computation difficult.
The computational tractability achieved in Koide et al. (2025) [5] relies on engineering-oriented assumptions motivated by numerical computation and simulation. Moreover, the applicability to real data is discussed in Koide et al. (2025) [4].
In contrast, the present study does not impose such engineering assumptions. Instead, by directly exploiting the probabilistic structure of stopping times for Lévy processes, we demonstrate that the expectation and variance of $T_{n}$ can be derived explicitly. Furthermore, we analyze the expectation of the inventory level, the expected total cost, and the asymptotic behavior of the total cost.
2. Setting
2.1. Fixed-Order-Quantity Policy under Intermittent Demand
Intermittent demand refers to a demand pattern in which demand does not occur continuously, but instead exhibits sudden arrivals after long periods of no demand. Such demand is characterized by strong uncertainty in both the timing of demand occurrences and the demand sizes. We model the cumulative demand under intermittent demand by a subordinator $\{D_{t}\}_{t\geq 0}$. In this model, the inventory level at time $t$ is given by the initial inventory minus the cumulative demand up to time $t$. Whenever the inventory level falls below a prescribed threshold, a replenishment of fixed size $Q$ is placed. Since the time at which the inventory reaches the threshold is stochastic, we define the corresponding stopping times. To model replenishment at each reorder point, the system is designed so that an order is placed whenever the cumulative demand exceeds a prescribed level. The order quantity is also incorporated explicitly. The formulation follows [5].
Definition 2.1 (First Passage Time of $\{D_{t}\}$).
Let $a>0$, $Q>0$, and $n\in\mathbb{N}$. The first passage time of the cumulative demand process $\{D_{t}\}$ is defined by
(1) $T_{n}^{a,Q}(=T_{n})=\inf\left\{s>0\,\middle|\,D_{s}\geq a+(n-1)Q\right\}.$
The parameter $a$ determines the reorder level, and $Q$ represents the order quantity. The random variable $T_{n}$ denotes the stochastic time of the $n$-th replenishment under the fixed-order-quantity policy. Using these quantities, we define the inventory control model under intermittent demand as follows.
Definition 2.2 (Inventory Level Process).
Let $x>0$ denote the initial inventory level. We consider a fixed-order-quantity policy with reorder point xax-a and order quantity $Q>0$. Let $\{D_{t}\}$ be the subordinator representing cumulative demand. The inventory level process $\{X_{t}\}$ is defined by
(2) $X_{t}=x-D_{t}+QR_{t},$
where
(3) $R_{t}=\sum_{n\geq 1}1_{\{T_{n}^{a,Q}<t\}}=\inf\{k\mid T_{k}^{a,Q}>t\}$
denotes the stochastic process representing the cumulative number of orders placed up to time $t$.
Taking expectations yields
$E[X_{t}]$ $=x-E[D_{t}]+QE[R_{t}]$
$=x-E[D_{t}]+Q\sum_{n\geq 1}E[1_{\{T_{n}<t\}}]\quad\text{(by Fubini's theorem)}$
Therefore,
(5) $E[X_{t}]$ $=x-E[D_{t}]+Q\sum_{n\geq 1}P(T_{n}<t).$
2.2. Expected Total Cost
In order to determine an optimal inventory control strategy, it is necessary to evaluate the associated costs explicitly. While Noba and Yamazaki [7] describe the total cost using a general functional form, here we decompose the total cost according to individual cost components.
Definition 2.3 (Expected Inventory Cost).
Let $C_{o},\ C_{h},\ C_{s.o.}>0$ denote the unit ordering cost, holding cost, and stockout cost, respectively. The total cost incurred up to time $t$ is defined by
(6) $C_{\mathrm{total}}(a,Q,t)=\underbrace{QC_{o}R_{t}}_{\text{ordering cost}}+\underbrace{C_{h}\int_{0}^{t}X_{s}ds}_{\text{holding cost}}+\underbrace{C_{s.o.}\int_{0}^{t}(-X_{s})^{+}ds}_{\text{stockout cost}}.$
The expectation $E[C_{\mathrm{total}}(a,Q,t)]$ is called the expected total cost.
Assuming that the stockout cost is negligible, the expected total cost becomes
$E[C_{\mathrm{total}}(a,Q,t)]$ $\simeq C_{o}QE[R_{t}]+C_{h}E\left[\int_{0}^{t}(x-D_{s}+QR_{s})ds\right]$
$=C_{o}QE[R_{t}]+C_{h}xt-C_{h}E\left[\int_{0}^{t}D_{s}ds\right]+C_{h}QE\left[\int_{0}^{t}R_{s}ds\right]$
$=C_{o}QE[R_{t}]+C_{h}xt-C_{h}\int_{0}^{t}E[D_{s}]ds+C_{h}QE\left[\int_{0}^{t}R_{s}ds\right].$
By Fubini’s theorem,
(7) $E[R_{t}]=E\!\left[\sum_{n\geq 1}\mathbf{1}_{\{T_{n}<t\}}\right]=\sum_{n\geq 1}P(T_{n}<t),$
and again by Fubini’s theorem,
(8) $E\!\left[\int_{0}^{t}R_{s}\,ds\right]=\int_{0}^{t}E[R_{s}]\,ds=\int_{0}^{t}\sum_{n\geq 1}P(T_{n}<s)\,ds.$
Hence,
$E[C_{\mathrm{total}}(a,Q,t)]$ $=C_{o}Q\sum_{n\geq 1}P(T_{n}<t)+C_{h}xt-C_{h}\int_{0}^{t}E[D_{s}]ds$
(9) $\quad+C_{h}Q\int_{0}^{t}\sum_{n\geq 1}P(T_{n}<s)\,ds.$
Remark 2.4.
For any $n\geq 1$, the identity
$\int_{0}^{t}\mathbf{1}_{\{T_{n}<s\}}\,ds=(t-T_{n})^{+}=t\,\mathbf{1}_{\{T_{n}<t\}}-T_{n}\mathbf{1}_{\{T_{n}<t\}}$
holds. Taking expectations and summing over $n$, and using Fubini’s theorem again due to non-negativity, we obtain
$E\!\left[\int_{0}^{t}R_{s}\,ds\right]$ $=\sum_{n\geq 1}E\!\left[\int_{0}^{t}\mathbf{1}_{\{T_{n}<s\}}\,ds\right]=\sum_{n\geq 1}E[(t-T_{n})^{+}]$
(10) $=\sum_{n\geq 1}\Bigl(t\,P(T_{n}<t)-E[T_{n}\mathbf{1}_{\{T_{n}<t\}}]\Bigr).$
Therefore, combining (10) and (8), we obtain
(11) $\sum_{n\geq 1}\Bigl(t\,P(T_{n}<t)-E[T_{n}\mathbf{1}_{\{T_{n}<t\}}]\Bigr)=\int_{0}^{t}\sum_{n\geq 1}P(T_{n}<s)\,ds.$
Regarding the Laplace transform of stopping times, the following well-known result holds.
Theorem (cf. J.Bertoin (1996) [1], Sato (1999) [8]).
Let $\{X_{t}\}$ be a subordinator (a one-dimensional, possibly killed, non-decreasing Lévy process), and define
$\tau_{b}^{+}=\inf\{t>0\mid X_{t}>b\}.$
Let $\Psi(\theta)=\log E[\exp(\theta X_{1})]$ denote the Laplace exponent. Then,
$E[\exp(-s\tau_{b}^{+})]=\exp(-b\Phi(s)),\quad\text{where }\Psi(\Phi(s))=s,\quad(s>0).$
Using this theorem, we will now examine the Laplace transforms of specific subordinators in detail.
3. Main Results
In (5) and (9), the terms involving $T_{n}$ are difficult to compute unless its distribution is explicitly known. In general, there are only few subordinators for which the distribution of the FPT can be written explicitly. In this section, we derive the Laplace transform of the FPT for a specific subordinator.
3.1. Case where the Cumulative Demand is a Drifted Poisson Process
Definition 3.1 (Cumulative Demand).
Let $\mu>0$, $\lambda>0$, and $\alpha>0$, and let $\{N_{t}\}_{t\geq 0}$ be a $\lambda$-Poisson process. We define the cumulative demand process $\{D_{t}\}$ as follows:
(12) $D_{t}=\mu t+\alpha N_{t}.$
We review the special function used in deriving the Laplace transform of the first passage time of $\{D_{t}\}$.
Definition 3.2 (Lambert W function [2]).
The complex function $W(z)$ satisfying
$W(z)e^{W(z)}=z$
is called the Lambert W function.
By choosing an appropriate branch, $W(z)$ can be regarded as a function from $\mathbb{R}$ to $\mathbb{R}$, and results concerning its derivative are also known.
Lemma 3.3 (Derivative of the Lambert W function[2]).
For the real-valued Lambert W function $W(x)$,
$\frac{d}{dx}W(x)=\frac{1}{(1+W(x))\exp(W(x))}=\frac{W(x)}{x(1+W(x))},\quad\text{if }x\not=0.$
The Laplace transform of the FPT for a drifted Poisson process can be expressed using the Lambert W function.
Theorem 3.4.
Let $T_{n}$ be the FPT of the drifted Poisson process $D_{t}=\mu t+\alpha N_{t}$ defined above. Let $W(x)$ denote the Lambert W function. Then the Laplace transform of $T_{n}$ is given as follows:
(13) $E[e^{-sT_{n}}]=\exp\left\{-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right\}.$
Proof.
We compute the Laplace exponent.
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\theta\mu+\lambda(e^{\theta\alpha}-1)$
The Laplace transform of $T_{n}$ is given by
$E[e^{-sT_{n}}]=\exp(-(a+(n-1)Q)\Phi(s))\quad\text{, where}\ \psi(\Phi(s))=s.$
We compute the inverse function of $\psi$.
$s$ $=\psi(\theta)=\theta\mu+\lambda(e^{\theta\alpha}-1)$
$s$ $=\theta\mu+\lambda(e^{\theta\alpha}-1)$
$\lambda e^{\theta\alpha}$ $=-\theta\mu+s+\lambda$
$\lambda e^{\theta\alpha}$ $=-\theta\mu+A\quad\quad(A=s+\lambda)$
$y$ $=\lambda\exp\left(-\alpha\frac{y-A}{\mu}\right)\quad\quad(y=-\theta\mu+A)$
$y$ $=\lambda\exp\left(\frac{\alpha A}{\mu}\right)\exp\left(-\frac{\alpha y}{\mu}\right)$
$y\exp\left(\frac{\alpha}{\mu}y\right)$ $=\lambda\exp\left(\frac{\alpha}{\mu}A\right)$
$\frac{\alpha}{\mu}y\exp\left(\frac{\alpha}{\mu}y\right)$ $=\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}A\right)$
$\frac{\alpha}{\mu}y$ $=W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}A\right)\right)\quad(W:\text{Lambert }W\text{ function})$
$-\alpha\theta+\frac{\alpha}{\mu}A$ $=W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}A\right)\right)$
$\theta$ $=-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}A\right)\right)+\frac{1}{\mu}A$
Therefore,
$\Phi(s)=-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right).$
Hence, the Laplace transform of $T_{n}$ is given by
$E[e^{-sT_{n}}]=\exp\left\{-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right\}.$
Corollary 3.5.
Under the assumptions of Theorem  3.4 , it holds that
(14) $E[T_{n}]$ $=\frac{a+(n-1)Q}{\mu+\alpha\lambda},$
(15) $\mathrm{Var}[T_{n}]$ $=\frac{\alpha^{2}\lambda(a+(n-1)Q)}{\left(\mu+\alpha\lambda\right)^{3}}.$
Proof.
Let
$\Phi(s)\coloneqq\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)-\frac{1}{\mu}\left(s+\lambda\right),\qquad K\coloneqq a+(n-1)Q$
be defined. We will organize the differentiation rules for the Laplace transform. We have
$\frac{d}{ds}E[e^{-sT_{n}}]$ $=\frac{d}{ds}\exp(-K\Phi(s))=-K\exp(-K\Phi(s))\frac{d\Phi}{ds}(s)$
$=-KE[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)=-KE[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)=-K\exp(-K\Phi(s))\frac{d\Phi}{ds}(s).$
We will also organize the second-order derivatives of the Laplace transform. We get
$\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]$ $=\frac{d}{ds}\left(-KE[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)\right)$
$=-K\left(\frac{d}{ds}E[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)+E[e^{-sT_{n}}]\frac{d^{2}\Phi}{ds^{2}}(s)\right)$
$=-K\left(-KE[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)\times\frac{d\Phi}{ds}(s)+E[e^{-sT_{n}}]\frac{d^{2}\Phi}{ds^{2}}(s)\right)$
$=-KE[e^{-sT_{n}}]\left(-K\left(\frac{d\Phi}{ds}(s)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(s)\right)$
$=-K\exp(-K\Phi(s))\left(-K\left(\frac{d\Phi}{ds}(s)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(s)\right).$
$\frac{d\Phi}{ds}(s)$ $=\frac{d}{ds}\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)$
$=-\frac{1}{\alpha}\frac{dW}{ds}\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}.$
Compute the first and second derivatives of $\Phi$. We have Let $B=\frac{\alpha}{\mu}$ and $C=\frac{\alpha\lambda}{\mu}$.
$\frac{d\Phi}{ds}(s)$ $=-\frac{1}{\alpha}\frac{dW}{ds}\left(C\exp\left(Bs+C\right)\right)+\frac{1}{\mu}.$
Furthermore,
$\frac{d^{2}\Phi}{ds^{2}}(s)$ $=-\frac{1}{\alpha}\frac{d^{2}W}{ds^{2}}\left(C\exp\left(Bs+C\right)\right).$
The derivative of $W(C\exp(Bs+C))$ can be computed as follows:
$\frac{dW}{ds}\left(C\exp\left(Bs+C\right)\right)$ $=\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}\left(C\exp\left(Bs+C\right)\right)^{-1}\left(BC\exp\left(Bs+C\right)\right)$
$=B\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}.$
Furthermore,
$\quad\frac{d^{2}W}{ds^{2}}\left(C\exp\left(Bs+C\right)\right)$
$=B\frac{d}{ds}\left(\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}\right)$
$=B\frac{dW}{ds}\left(C\exp\left(Bs+C\right)\right)\frac{1}{1+W\left(C\exp\left(Bs+C\right)\right)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+B\times W\left(C\exp\left(Bs+C\right)\right)\frac{d}{ds}\left(\frac{1}{1+W\left(C\exp\left(Bs+C\right)\right)}\right)$
$=B\times B\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}\times\frac{1}{1+W\left(C\exp\left(Bs+C\right)\right)}$
$\quad\quad\quad\quad\quad\quad\quad\quad+B\times W\left(C\exp\left(Bs+C\right)\right)\times(-1)\frac{\frac{dW}{ds}\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{2}}$
$=B\times B\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}\times\frac{1}{1+W\left(C\exp\left(Bs+C\right)\right)}$
$\quad\quad-B\times W\left(C\exp\left(Bs+C\right)\right)\times\frac{1}{(1+W\left(C\exp\left(Bs+C\right)\right))^{2}}\times B\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}$
$=B^{2}\frac{W\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{2}}-B^{2}\frac{(W\left(C\exp\left(Bs+C\right)\right))^{2}}{(1+W\left(C\exp\left(Bs+C\right)\right))^{3}}$
$=B^{2}\frac{W\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{2}}\left(1-\frac{(W\left(C\exp\left(Bs+C\right)\right))}{(1+W\left(C\exp\left(Bs+C\right)\right))}\right)$
$=B^{2}\frac{W\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{2}}\times\frac{1}{1+W\left(C\exp\left(Bs+C\right)\right)}$
$=B^{2}\frac{W\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{3}}.$
It follows that
$\quad\frac{d}{ds}E[e^{-sT_{n}}]$
$=-K\exp(-K\Phi(s))\frac{d\Phi}{ds}(s)$
$=-(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right]$
$\quad\times\left(-\frac{1}{\alpha}B\frac{W\left(C\exp\left(Bs+C\right)\right)}{1+W\left(C\exp\left(Bs+C\right)\right)}+\frac{1}{\mu}\right)$
$=-(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}\left(s+\lambda\right)\right)\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right]$
$\quad\times\left(-\frac{1}{\mu}\frac{W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)}{1+W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)}+\frac{1}{\mu}\right).$
Therefore, the expected value of $E[T_{n}]$ is given by:
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}$
$=(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)-\frac{\lambda}{\mu}\right)\right]$
$\quad\times\left(-\frac{1}{\mu}\frac{W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)}{1+W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)}+\frac{1}{\mu}\right)$
$=(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(\frac{1}{\alpha}\frac{\alpha\lambda}{\mu}-\frac{\lambda}{\mu}\right)\right]\times\left(-\frac{1}{\mu}\frac{\frac{\alpha\lambda}{\mu}}{1+\frac{\alpha\lambda}{\mu}}+\frac{1}{\mu}\right)$
$=\frac{a+(n-1)Q}{\mu+\alpha\lambda}.$
Similarly, calculating the second moment yields:
$\quad\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]$
$=\frac{d}{ds}\left(-KE[e^{-sT_{n}}]\frac{d\Phi}{ds}(s)\right)$
$=-K\exp(-K\Phi(s))\left(-K\left(\frac{d\Phi}{ds}(s)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(s)\right)$
$=-(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha}{\mu}\left(s+\lambda\right)}\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right]$
$\quad\times\left(-(a+(n-1)Q)\left(-\frac{1}{\alpha}\frac{dW}{ds}\left(C\exp\left(Bs+C\right)\right)+\frac{1}{\mu}\right)^{2}\right$.
$\quad\left.-\frac{1}{\alpha}B^{2}\frac{W\left(C\exp\left(Bs+C\right)\right)}{(1+W\left(C\exp\left(Bs+C\right)\right))^{3}}\right)$
$=-(a+(n-1)Q)\exp\left[-(a+(n-1)Q)\left(-\frac{1}{\alpha}W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha}{\mu}\left(s+\lambda\right)}\right)+\frac{1}{\mu}\left(s+\lambda\right)\right)\right]$
$\quad\times\left(-(a+(n-1)Q)\left(-\frac{1}{\mu}\frac{W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)}{1+W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)}+\frac{1}{\mu}\right)^{2}\right$.
$\quad\left.-\frac{\alpha}{\mu^{2}}\frac{W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)}{\left(1+W\left(\frac{\alpha\lambda}{\mu}\exp\left(\frac{\alpha}{\mu}s+\frac{\alpha\lambda}{\mu}\right)\right)\right)^{3}}\right).$
Hense,
$E[T_{n}^{2}]$ $=-(a+(n-1)Q)\times\left(-(a+(n-1)Q)\left(-\frac{1}{\mu}\frac{W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)}{1+W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)}+\frac{1}{\mu}\right)^{2}\right$.
$\quad\left.-\frac{\alpha}{\mu^{2}}\frac{W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)}{\left(1+W\left(\frac{\alpha\lambda}{\mu}e^{\frac{\alpha\lambda}{\mu}}\right)\right)^{3}}\right)$
$=-(a+(n-1)Q)\times\left(-(a+(n-1)Q)\left(-\frac{1}{\mu}\frac{\frac{\alpha\lambda}{\mu}}{1+\frac{\alpha\lambda}{\mu}}+\frac{1}{\mu}\right)^{2}\right$.
$\quad\left.-\frac{\alpha}{\mu^{2}}\frac{\frac{\alpha\lambda}{\mu}}{\left(1+\frac{\alpha\lambda}{\mu}\right)^{3}}\right)$
$=\left(\frac{a+(n-1)Q}{\mu+\alpha\lambda}\right)^{2}+\frac{\alpha^{2}\lambda(a+(n-1)Q)}{\left(\mu+\alpha\lambda\right)^{3}}.$
From these results, we obtain
$\text{Var}[T_{n}]=E[T_{n}^{2}]-E[T_{n}]^{2}=\frac{\alpha^{2}\lambda(a+(n-1)Q)}{\left(\mu+\alpha\lambda\right)^{3}}.$
3.2. Case where the Cumulative Demand is a Drifted Compound Poisson Process
As a generalization of the drifted Poisson process, considering a drifted compound Poisson process is not only a mathematical generalization but also a valuable extension in practice, since it can reproduce situations in which the magnitude of sudden demand changes is random each time.
Definition 3.6 (Cumulative Demand).
Let $\mu>0$, $\lambda>0$, and let $\{N_{t}\}_{t\geq 0}$ be a $\lambda$-Poisson process. Let $\{J_{k}\}_{n\geq 1}$ be a sequence of random variables that are independent of $\{D_{t}\}_{t\geq 0}$ and independent and identically distributed. We define the cumulative demand process $\{D_{t}\}_{t\geq 0}$ as follows:
(16) $D_{t}=\mu t+\sum_{k=1}^{N_{t}}J_{k}.$
In this case, let $T_{n}$ be the FPT defined above, and let $\psi(\theta)$ be the Laplace exponent of $D_{t}$. Then,
$E[e^{-sT_{n}}]=\exp(-(a+(n-1)Q)\Phi(s))\quad\text{,where}\ \psi(\Phi(s))=s.$
Let $L_{J_{1}}(\theta)$ denote the Laplace transform of $J_{1}$, and compute the Laplace exponent $\psi(\theta)$.
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\log E\left[\exp\left(\theta\mu+\theta\sum_{k=1}^{N_{1}}J_{k}\right)\right]=\theta\mu+\log E\left[\exp\left(\theta\sum_{k=1}^{N_{1}}J_{k}\right)\right]$
$E\left[\exp\left(\theta\sum_{k=1}^{N_{1}}J_{k}\right)\right]$ $=E\left[E\left[\exp\left(\theta\sum_{k=1}^{N_{1}}J_{k}\right)\ \middle|\ N_{1}\right]\right]$
$=\sum_{m\geq 0}E\left[\exp\left(\theta\sum_{k=1}^{m}J_{k}\right)\right]P(N_{1}=m)$
$=\sum_{m\geq 0}E\left[\exp(\theta J_{1})\right]^{m}P(N_{1}=m)\quad\because\ J_{k}\ \text{i.i.d.}$
$=\sum_{m\geq 0}{L_{J_{1}}(\theta)}^{m}e^{-\lambda}\frac{\lambda^{m}}{m!}$
$=e^{-\lambda}\exp(\lambda L_{J_{1}}(\theta))$
Therefore,
$\log E\left[\exp\left(\theta\sum_{k=1}^{N_{1}}J_{k}\right)\right]$ $=\log\exp(\lambda(L_{J_{1}}(\theta)-1))$
$=\lambda(L_{J_{1}}(\theta)-1).$
Hence,
(17) $\psi(\theta)=\theta\mu+\lambda(L_{J_{1}}(\theta)-1).$
Case where the Jump Sizes Follow an Exponential Distribution
Theorem 3.7.
Let $\{D_{t}\}_{t\geq 0}$ be defined as in (16), and assume that $J_{k}$ are independent and identically distributed and follow an exponential distribution with parameter $\eta$. In this case, for $\theta<\eta$,
(18) $E[e^{-sT_{n}}]=\exp\left(-(a+(n-1)Q)\frac{1}{2\mu}\left\{(\mu\eta+s+\lambda)-\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}\right\}\right).$
Proof.
Since $L_{J_{1}}(\theta)=\frac{\eta}{\eta-\theta}$, it follows from (17) that
$\psi(\theta)=\theta\mu+\lambda\left(\frac{\eta}{\eta-\theta}-1\right)=\theta\mu+\frac{\lambda\theta}{\eta-\theta}.$
We compute the inverse function of $\psi$.
$s=\psi(\theta)$ $=\theta\mu+\lambda\left(\frac{\eta}{\eta-\theta}-1\right)=\theta\mu+\frac{\lambda\theta}{\eta-\theta}$
$\theta\mu(\eta-\theta)+\lambda\theta$ $=s(\eta-\theta)$
$\mu\theta^{2}-(\mu\eta+s+\lambda)\theta+s\eta$ $=0$
$\Phi(s)$ $=\theta=\frac{1}{2\mu}\left\{(\mu\eta+s+\lambda)-\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}\right\}.$
The other solution of the quadratic equation does not satisfy $\theta<\eta$, and is therefore not appropriate. Hence,
$E[e^{-sT_{n}}]$ $=\exp(-(a+(n-1)Q)\Phi(s))$
$=\exp\left(-(a+(n-1)Q)\frac{1}{2\mu}\left\{(\mu\eta+s+\lambda)-\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}\right\}\right).$
Corollary 3.8.
Under the assumptions of Theorem  3.7, it holds that
(19) $E[T_{n}]=\frac{\eta(a+(n-1)Q)}{\mu\eta+\lambda},$
(20) $\text{Var}[T_{n}]=\frac{2\eta\lambda(a+(n-1)Q)}{(\mu\eta+\lambda)^{3}}.$
Proof.
The calculation follows the same approach as Corollary 3.5. Let
$\Phi(s)\coloneqq\frac{1}{2\mu}\left\{(\mu\eta+s+\lambda)-\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}\right\},\quad K\coloneqq a+(n-1)Q$
Recall the following result regarding the differentiation of the Laplace transform of $T_{n}$:
$\frac{d}{ds}E[e^{-sT_{n}}]$ $=-K\exp(-K\Phi(s))\frac{d\Phi}{ds}(s).$
$\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]=-K\exp(-K\Phi(s))\left(-K\left(\frac{d\Phi}{ds}(s)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(s)\right).$
Compute the first and second derivatives of $\Phi$.
$\frac{d\Phi}{ds}(s)$ $=\frac{1}{2\mu}\left\{1-\frac{2(\mu\eta+s+\lambda)-4\mu\eta}{2\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}$
$=\frac{1}{2\mu}\left\{1-\frac{2s+2\lambda-2\mu\eta}{2\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}$
$=\frac{1}{2\mu}\left\{1-\frac{s+\lambda-\mu\eta}{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}.$
$\frac{d^{2}\Phi}{ds^{2}}(s)$ $=-\frac{1}{2\mu}\times\frac{\left\{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}-(s+\lambda-\mu\eta)\frac{(s+\lambda-\mu\eta)}{\sqrt{(\mu\eta+s+\lambda)^{2}-4\mu s\eta}}\right\}}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}$
$=-\frac{1}{2\mu}\times\frac{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta-(s+\lambda-\mu\eta)^{2}\right\}}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}$
$=-\frac{1}{2\mu}\times\frac{\left\{4\mu\eta(s+\lambda)-4\mu s\eta\right\}}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}$
$=-\frac{2\eta(s+\lambda)-2\mu s\eta}{\left\{(\mu\eta+s+\lambda)^{2}-4\mu s\eta\right\}^{\frac{3}{2}}}.$
Moreover,We verify the value at 0.
$\Phi(0)=0.$
$\frac{d\Phi}{ds}(0)$ $=\frac{1}{2\mu}\left\{1-\frac{\lambda-\mu\eta}{\mu\eta+\lambda}\right\}$
$=\frac{1}{2\mu}\left\{\frac{2\mu\eta}{\mu\eta+\lambda}\right\}=\frac{\eta}{\mu\eta+\lambda}.$
$\frac{d^{2}\Phi}{ds^{2}}(0)=-\frac{2\eta\lambda}{\left\{(\mu\eta+\lambda)^{2}\right\}^{\frac{3}{2}}}=-\frac{2\eta\lambda}{(\mu\eta+\lambda)^{3}}.$
Therefore, the expectation is given by
$E[T_{n}]=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}=K\exp(-K\Phi(0))\frac{d\Phi}{ds}(0)=\frac{\eta(a+(n-1)Q)}{\mu\eta+\lambda}.$
Regarding the second moment, we have
$E[T_{n}^{2}]$ $=\left.\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]\right|_{s=0}$
$=-K\exp(-K\Phi(0))\left(-K\left(\frac{d\Phi}{ds}(0)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(0)\right)$
$=(a+(n-1)Q)^{2}\left(\frac{\eta}{\mu\eta+\lambda}\right)^{2}+(a+(n-1)Q)\frac{2\eta\lambda}{(\mu\eta+\lambda)^{3}}$
$=\frac{\eta^{2}(a+(n-1)Q)^{2}}{(\mu\eta+\lambda)^{2}}+\frac{2\eta\lambda(a+(n-1)Q)}{(\mu\eta+\lambda)^{3}}.$
As for the variance, we obtain
$\text{Var}[T_{n}]=E[T_{n}^{2}]-E[T_{n}]^{2}=\frac{2\eta\lambda(a+(n-1)Q)}{(\mu\eta+\lambda)^{3}}.$
Case where the Jump Sizes Follow a Gamma Distribution
As a generalization of Theorem 3.7, we consider a compound Poisson process whose jump sizes follow a Gamma distribution. In this case, the Laplace transform cannot be expressed in closed form. However, since its derivatives at the origin can be computed explicitly, the expectation and variance can still be obtained.
Theorem 3.9.
Let $\{D_{t}\}_{t\geq 0}$ be defined as in (16), and assume that $J_{k}$ are independent and identically distributed and follow a Gamma distribution with parameters $\beta$ and $\eta$. Then,
$E[T_{n}]$ $=\frac{\eta(a+(n-1)Q)}{\mu\eta+\beta\lambda},$
$\mathrm{Var}[T_{n}]$ $=\frac{\beta(\beta+1)\eta\lambda(a+(n-1)Q)}{(\mu\eta+\beta\lambda)^{3}}.$
Proof.
Let $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$. Then, for $\theta<\eta$, the Laplace Exponent $\Psi$ can be calculated as follows:
$\psi(\theta)=\theta\mu+\lambda\left(\left(\frac{\eta}{\eta-\theta}\right)^{\beta}-1\right).$
$\frac{d\psi}{d\theta}(\theta)=\mu+\lambda\beta\eta^{\beta}(\eta-\theta)^{-(\beta+1)}.$
$\frac{d^{2}\psi}{d\theta^{2}}(\theta)=\lambda\beta(\beta+1)\eta^{\beta}(\eta-\theta)^{-(\beta+2)}.$
Let $\psi^{-1}=\Phi$ denote the inverse function of $\psi$. Then, it holds that
$\Phi(0)=0\quad(\because\psi(0)=0),$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\lambda\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\beta\lambda},$
$\frac{d^{2}\Phi}{ds^{2}}(0)=-\frac{\cfrac{d^{2}\psi}{d\theta^{2}}(0)}{\left(\cfrac{d\psi}{d\theta}(0)\right)^{3}}=-\frac{\lambda\beta(\beta+1)\eta^{-2}}{(\mu+\lambda\beta\eta^{-1})^{3}}=-\frac{\beta(\beta+1)\eta\lambda}{(\mu\eta+\beta\lambda)^{3}}.$
Therefore, the expectation is given by
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}=K\exp(-K\Phi(0))=\frac{d\Phi}{ds}(0)\frac{\eta(a+(n-1)Q)}{\mu\eta+\beta\lambda}.$
The second moment is given by
$E[T_{n}^{2}]$ $=\left.\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]\right|_{s=0}$
$=-K\exp(-K\Phi(0))\left(-K\left(\frac{d\Phi}{ds}(0)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(0)\right)$
$=-(a+(n-1)Q)\left(-(a+(n-1)Q)\left(\frac{\eta}{\mu\eta+\beta\lambda}\right)^{2}-\frac{\beta(\beta+1)\eta\lambda}{(\mu\eta+\beta\lambda)^{3}}\right)$
$=\left(\frac{\eta(a+(n-1)Q)}{\mu\eta+\beta\lambda}\right)^{2}+\frac{\beta(\beta+1)\eta\lambda(a+(n-1)Q)}{(\mu\eta+\beta\lambda)^{3}}.$
The variance is given by
$\text{Var}[T_{n}]=E[T_{n}^{2}]-E[T_{n}]^{2}=\frac{\beta(\beta+1)\eta\lambda(a+(n-1)Q)}{(\mu\eta+\beta\lambda)^{3}}.$
3.3. Generalization of the Cumulative Demand
Based on the previous discussion, we further generalize the model.
Definition 3.10 (Generalized Cumulative Demand).
Let $\mu,\alpha,\lambda,\lambda^{\prime}>0$. Let $\{N_{t}\}_{t\geq 0}$ be a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ be a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ be a sequence of independent and identically distributed random variables. Assume that $\{N_{t}\}_{t\geq 0}$, $\{N^{\prime}_{t}\}_{t\geq 0}$, and $\{J_{k}\}_{k\geq 1}$ are mutually independent. We define the cumulative demand process $\{D_{t}\}_{t\geq 0}$ as
(21) $D_{t}=\mu t+\alpha N_{t}+\sum_{k=1}^{N^{\prime}_{t}}J_{k}.$
In this case, for
$T_{n}=\inf\left\{s>T_{n-1}\,\middle|\,D_{s}\geq a+(n-1)Q\right\},$
we derive $E[e^{-sT_{n}}]$, $E[T_{n}]$, and $\text{Var}[T_{n}]$. Let $\psi(\theta)$ denote the Laplace exponent of $D_{t}$. Then,
$E[e^{-sT_{n}}]=\exp(-(a+(n-1)Q)\Phi(s))\quad\text{,where}\ \psi(\Phi(s))=s.$
Let $L_{J_{1}}(\theta)$ denote the Laplace transform of $J_{1}$. We compute the Laplace exponent $\psi(\theta)$:
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\theta\mu+\lambda(e^{\theta\alpha}-1)+\lambda^{\prime}(L_{J_{1}}(\theta)-1).$
Theorem 3.11.
Let $\{D_{t}\}_{t\geq 0}$ be defined as in (21), and assume that $J_{k}$ are independent and identically distributed and follow a Gamma distribution with parameters $\beta,\eta$. Then,
(22) $E[T_{n}]=\frac{\eta(a+(n-1)Q)}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}},$
(23) $\text{Var}[T_{n}]=\frac{(\alpha^{2}\eta^{3}\lambda+\beta(\beta+1)\eta\lambda^{\prime})(a+(n-1)Q)}{(\mu\eta+\alpha\eta\lambda+\beta\lambda^{\prime})^{3}}.$
Proof.
Since $L_{J_{1}}=\left(\frac{\eta}{\eta-\theta}\right)^{\beta}$, for $\theta<\eta$, the Laplace Exponent $\Psi$ can be calculated as follows:
$\psi(\theta)=\log E[e^{\theta D_{1}}]=\theta\mu+\lambda(e^{\theta\alpha}-1)+\lambda^{\prime}\left(\left(\frac{\eta}{\eta-\theta}\right)^{\beta}-1\right),$
$\frac{d\psi}{d\theta}(\theta)=\mu+\alpha\lambda e^{\alpha\theta}+\lambda^{\prime}\beta\eta^{\beta}(\eta-\theta)^{-(\beta+1)},$
$\frac{d^{2}\psi}{d\theta^{2}}(\theta)=\alpha^{2}\lambda e^{\alpha\theta}+\lambda^{\prime}\beta(\beta+1)\eta^{\beta}(\eta-\theta)^{-(\beta+2)}.$
Let $\Phi$ denote the inverse function of $\psi$. Then, it holds that
$\Phi(0)=0\quad(\because\psi(0)=0),$
$\frac{d\Phi}{ds}(0)=\frac{1}{\cfrac{d\psi}{d\theta}(0)}=\frac{1}{\mu+\alpha\lambda+\lambda^{\prime}\beta\eta^{-1}}=\frac{\eta}{\mu\eta+\alpha\lambda\eta+\beta\lambda^{\prime}},$
$\frac{d^{2}\Phi}{ds^{2}}(0)=-\frac{\cfrac{d^{2}\psi}{d\theta^{2}}(0)}{\left(\cfrac{d\psi}{d\theta}(0)\right)^{3}}=-\frac{\alpha^{2}\lambda+\lambda^{\prime}\beta(\beta+1)\eta^{-2}}{(\mu+\alpha\lambda+\lambda^{\prime}\beta\eta^{-1})^{3}}=-\frac{\alpha^{2}\eta^{3}\lambda+\beta(\beta+1)\eta\lambda^{\prime}}{(\mu\eta+\alpha\lambda\eta+\beta\lambda^{\prime})^{3}}.$
Therefore, we have
$E[T_{n}]$ $=-\left.\frac{d}{ds}E[e^{-sT_{n}}]\right|_{s=0}=K\exp(-K\Phi(0))\frac{d\Phi}{ds}(0)=\frac{\eta(a+(n-1)Q)}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}},$
$E[T_{n}^{2}]$ $=\left.\frac{d^{2}}{ds^{2}}E[e^{-sT_{n}}]\right|_{s=0}$
$=-K\exp(-K\Phi(0))\left(-K\left(\frac{d\Phi}{ds}(0)\right)^{2}+\frac{d^{2}\Phi}{ds^{2}}(0)\right)$
$=-(a+(n-1)Q)\left(-(a+(n-1)Q)\left(\frac{\eta}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}}\right)^{2}\right$.
$\quad-\left.\frac{(\alpha^{2}\eta^{3}\lambda+\beta(\beta+1)\eta\lambda^{\prime})}{(\mu\eta+\alpha\eta\lambda+\beta\lambda^{\prime})^{3}}\right)$
$=\left(\frac{\eta(a+(n-1)Q)}{\mu\eta+\alpha\lambda+\beta\lambda^{\prime}}\right)^{2}+\frac{(\alpha^{2}\eta^{3}\lambda+\beta(\beta+1)\eta\lambda^{\prime})(a+(n-1)Q)}{(\mu\eta+\alpha\eta\lambda+\beta\lambda^{\prime})^{3}}.$
For the variance, we obtain
$\text{Var}[T_{n}]=E[T_{n}^{2}]-E[T_{n}]^{2}=\frac{(\alpha^{2}\eta^{3}\lambda+\beta(\beta+1)\eta\lambda^{\prime})(a+(n-1)Q)}{(\mu\eta+\alpha\eta\lambda+\beta\lambda^{\prime})^{3}}.$
3.4. On the Expected Total Cost
Although we do not directly use the distribution of $T_{n}$, we consider the generalized cumulative demand process and investigate the expected value of the total inventory cost derived from it. Consider the cumulative demand process
$D_{t}=\mu t+\alpha N_{t}+\sum_{k=1}^{N^{\prime}_{t}}J_{k}.$
Here, $\{N_{t}\}_{t\geq 0}$ is a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ is a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ is a sequence of nonnegative independent and identically distributed random variables. We assume that these processes are mutually independent.
Lemma 3.12.
For any $s>0$ and $b>0$,
(24) $P(T_{n}<s)=P(D_{s}\geq b)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}p_{i}(s)\,q_{j}(s)\,P\!\left(\sum_{k=1}^{j}J_{k}\geq b-\mu s-\alpha i\right)$
holds, where
$p_{i}(s)=e^{-\lambda s}\frac{(\lambda s)^{i}}{i!},\qquad q_{j}(s)=e^{-\lambda^{\prime}s}\frac{(\lambda^{\prime}s)^{j}}{j!}.$
Proof.
Since the cumulative demand process $D_{t}$ is nondecreasing, $\{T_{n}<s\}\iff\{D_{s}\geq b\}$ holds, and thus $P(T_{n}<s)=P(D_{s}\geq b)$.
$P(D_{s}\geq b)=\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}P(D_{s}\geq b\mid N_{s}=i,N^{\prime}_{s}=j)\,P(N_{s}=i,N^{\prime}_{s}=j).$
Because the Poisson processes $N_{s}$ and $N^{\prime}_{s}$ are independent,
$P(N_{s}=i,N^{\prime}_{s}=j)=P(N_{s}=i)P(N^{\prime}_{s}=j)=p_{i}(s)\,q_{j}(s).$
Under the condition $(N_{s},N^{\prime}_{s})=(i,j)$,
$D_{s}=\mu s+\alpha i+\sum_{k=1}^{j}J_{k}\quad\text{a.s.}$
Therefore,
$P(D_{s}\geq b\mid N_{s}=i,N^{\prime}_{s}=j)=P\!\left(\sum_{k=1}^{j}J_{k}\geq b-\mu s-\alpha i\right).$
Lemma 3.13.
Let $\{J_{k}\}_{k\geq 1}$ be independent and identically distributed random variables following a Gamma distribution with parameters $\beta,\eta>0$. Then for any integer $j\geq 1$ and $x\in\mathbb{R}$,
(25) $P\!\left(\sum_{k=1}^{j}J_{k}\geq x\right)=\begin{cases}1,&x\leq 0,\\[6.0pt] \dfrac{\Gamma(j\beta,\eta x)}{\Gamma(j\beta)},&x>0,\end{cases}$
where $\Gamma(\cdot,\cdot)$ denotes the upper incomplete gamma function.
In particular, when $\beta=1$,
(26) $P\!\left(\sum_{k=1}^{j}J_{k}\geq x\right)=e^{-\eta x}\sum_{r=0}^{j-1}\frac{(\eta x)^{r}}{r!}.$
Proof.
If $x\leq 0$, since $J_{k}\geq 0$, it is clear that $\sum_{k=1}^{j}J_{k}\geq x$.
Assume $x>0$. Using the Laplace transform of the Gamma distribution
$E[e^{sJ_{1}}]=\left(\frac{\eta}{\eta-s}\right)^{\beta}$
and independence, we obtain
$E\!\left[e^{-s\sum_{k=1}^{j}J_{k}}\right]=\left(\frac{\eta}{\eta-s}\right)^{j\beta}.$
This coincides with the Laplace transform of the Gamma distribution with parameters $j\beta,\eta$. Therefore,
$\sum_{k=1}^{j}J_{k}$
follows a Gamma distribution with parameters $j\beta,\eta$. Hence,
$P\!\left(\sum_{k=1}^{j}J_{k}\geq x\right)=\int_{x}^{\infty}\frac{\eta^{j\beta}}{\Gamma(j\beta)}u^{j\beta-1}e^{-\eta u}\,du.$
By the change of variables $y=\eta u$,
$P\!\left(\sum_{k=1}^{j}J_{k}\geq x\right)=\frac{1}{\Gamma(j\beta)}\int_{\eta x}^{\infty}y^{j\beta-1}e^{-y}\,dy=\frac{\Gamma(j\beta,\eta x)}{\Gamma(j\beta)}.$
When $\beta=1$, $j\beta=j$ is an integer. Consider
$\Gamma(j,z)=\int_{z}^{\infty}y^{j-1}e^{-y}\,dy.$
Let $I_{j-1}(z):=\int_{z}^{\infty}y^{j-1}e^{-y}\,dy$.
$I_{j-1}(z)=\Bigl[-y^{j-1}e^{-y}\Bigr]_{z}^{\infty}+(j-1)\int_{z}^{\infty}y^{j-2}e^{-y}\,dy.$
Since $\lim_{y\to\infty}y^{j-1}e^{-y}=0$,
$I_{j-1}(z)=z^{j-1}e^{-z}+(j-1)I_{j-2}(z).$
Repeating integration by parts recursively yields
$I_{j-1}(z)=e^{-z}\Bigl(z^{j-1}+(j-1)z^{j-2}+(j-1)(j-2)z^{j-3}+\cdots+(j-1)!\Bigr).$
Rearranging coefficients gives
$I_{j-1}(z)=(j-1)!e^{-z}\sum_{r=0}^{j-1}\frac{z^{r}}{r!}.$
Thus,
$\Gamma(j,z)=(j-1)!e^{-z}\sum_{r=0}^{j-1}\frac{z^{r}}{r!}.$
Therefore,
$\Gamma(j,\eta x)=(j-1)!e^{-\eta x}\sum_{r=0}^{j-1}\frac{(\eta x)^{r}}{r!}.$
Remark 3.14.
The same result can also be obtained by performing a direct computation in the case where the jump sizes $J_{k}$ follow an exponential distribution.
Theorem 3.15.
Consider the cumulative demand process
$D_{t}=\mu t+\alpha N_{t}+\sum_{k=1}^{N^{\prime}_{t}}J_{k}.$
Here, $\{N_{t}\}_{t\geq 0}$ is a $\lambda$-Poisson process, $\{N^{\prime}_{t}\}_{t\geq 0}$ is a $\lambda^{\prime}$-Poisson process, and $\{J_{k}\}_{k\geq 1}$ is a sequence of nonnegative independent and identically distributed random variables. Assume that these are mutually independent. Then, for any $s>0$ and $b>0$,
$E[C_{\mathrm{total}}(a,Q,t)]$ $=C_{o}Q\sum_{n\geq 1}P(D_{t}\geq a+(n-1)Q)$
$\quad+C_{h}xt-\frac{C_{h}t^{2}}{2}\bigl(\mu+\alpha\lambda+\lambda^{\prime}E[J_{1}]\bigr)$
(27) $\quad+C_{h}Q\int_{0}^{t}\sum_{n\geq 1}P(D_{s}\geq a+(n-1)Q)\,ds.$
Proof.
The result follows from (9) and (24). ∎
If the distribution of $J_{k}$ is specified explicitly, we obtain the following corollary.
Corollary 3.16.
Let $\{J_{k}\}_{k\geq 1}$ be independent and identically distributed random variables following a Gamma distribution with parameters $\beta,\eta>0$. Then the expected total cost is given by
$\quad E[C_{\mathrm{total}}(a,Q,t)]$
$=C_{o}Q\sum_{n\geq 1}\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}p_{i}(t)q_{j}(t)\,P\!\left(\sum_{k=1}^{j}J_{k}\geq a+(n-1)Q-\mu t-\alpha i\right)$
$\quad+C_{h}xt-\frac{C_{h}t^{2}}{2}\Bigl(\mu+\alpha\lambda+\lambda^{\prime}\tfrac{\beta}{\eta}\Bigr)$
(28) $\quad+C_{h}Q\int_{0}^{t}\sum_{n\geq 1}\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}p_{i}(s)q_{j}(s)\,P\!\left(\sum_{k=1}^{j}J_{k}\geq a+(n-1)Q-\mu s-\alpha i\right)ds,$
where
$p_{i}(s)=e^{-\lambda s}\frac{(\lambda s)^{i}}{i!},\qquad q_{j}(s)=e^{-\lambda^{\prime}s}\frac{(\lambda^{\prime}s)^{j}}{j!},$
and
$P\!\left(\sum_{k=1}^{j}J_{k}\geq y\right)=\begin{cases}1,&y\leq 0,\\[6.0pt] \dfrac{\Gamma(j\beta,\eta y)}{\Gamma(j\beta)},&y>0,\end{cases}$
holds.In particular, when $\beta=1$,
$E[C_{\mathrm{total}}(a,Q,t)]$ $=C_{o}Q\sum_{n\geq 1}\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}p_{i}(t)q_{j}(t)\,\Phi_{j}(a+(n-1)Q-\mu t-\alpha i)$
$\quad+C_{h}xt-\frac{C_{h}t^{2}}{2}\Bigl(\mu+\alpha\lambda+\tfrac{\lambda^{\prime}}{\eta}\Bigr)$
(29) $\quad+C_{h}Q\int_{0}^{t}\sum_{n\geq 1}\sum_{i=0}^{\infty}\sum_{j=0}^{\infty}p_{i}(s)q_{j}(s)\,\Phi_{j}(a+(n-1)Q-\mu s-\alpha i)ds.$
Here,
$\phi_{j}(y)=P\!\left(\sum_{k=1}^{j}J_{k}\geq y\right)=\begin{cases}1,&y\leq 0,\\[6.0pt] \displaystyle e^{-\eta y}\sum_{r=0}^{j-1}\frac{(\eta y)^{r}}{r!},&y>0.\end{cases}$
Theorem 3.17 (Comparison between Finite-Horizon and Long-Run Average).
Assume $E[J_{1}]<\infty$ and define
$m\coloneq\bigl(\mu+\alpha\lambda+\lambda^{\prime}E[J_{1}]\bigr).$
Then,
(30) $\lim_{t\to\infty}\frac{1}{t}E[C_{\mathrm{total}}(a,Q,t)]=C_{o}m+C_{h}x,$
and this limit does not depend on $Q$ (nor on $a$).
Proof.
By monotonicity,
$\{T_{n}<t\}\iff\{D_{t}\geq a+(n-1)Q\}$
holds, and therefore
$E[R_{t}]=\sum_{n\geq 1}P(T_{n}<t)=\sum_{n\geq 1}P(D_{t}\geq a+(n-1)Q).$
Hence, from the previously derived representation of the total cost,
$E[C_{\mathrm{total}}(a,Q,t)]=C_{o}Q\,E[R_{t}]+C_{h}xt-C_{h}\int_{0}^{t}E[D_{s}]\,ds+C_{h}Q\int_{0}^{t}E[R_{s}]\,ds.$
Next, we compute $E[D_{s}]$. By independence and linearity,
$E[D_{s}]=\mu s+\alpha\,E[N_{s}]+E[N^{\prime}_{s}]E[J_{1}]=\bigl(\mu+\alpha\lambda+\lambda^{\prime}E[J_{1}]\bigr)s=:ms.$
Thus,
$\int_{0}^{t}E[D_{s}]\,ds=\frac{m}{2}t^{2}.$
To establish the long-run average, we provide upper and lower bounds relating $R_{t}$ and $D_{t}$. Since $R_{t}$ counts the number of thresholds reached by $D_{t}$, we have
$a+(R_{t}-1)Q\leq D_{t}<a+R_{t}Q.$
Therefore,
$\frac{D_{t}-a}{Q}\leq R_{t}\leq\frac{D_{t}-a}{Q}+1.$
Taking expectations and dividing by $t$,
$\frac{1}{t}\frac{E[D_{t}]-a}{Q}\leq\frac{E[R_{t}]}{t}\leq\frac{1}{t}\frac{E[D_{t}]-a}{Q}+\frac{1}{t}.$
Since $E[D_{t}]=mt$, letting $t\to\infty$ yields
$\lim_{t\to\infty}\frac{E[R_{t}]}{t}=\frac{m}{Q}.$
Furthermore,
$\frac{1}{t^{2}}\int_{0}^{t}E[R_{s}]\,ds=\int_{0}^{1}\frac{E[R_{ut}]}{t}\,du\longrightarrow\int_{0}^{1}\frac{m}{Q}u\,du=\frac{m}{2Q}.$
Hence,
$\int_{0}^{t}E[R_{s}]\,ds\sim\frac{m}{2Q}t^{2}.$
Substituting these into the total cost expression,
$\frac{1}{t}E[C_{\mathrm{total}}(a,Q,t)]$ $=C_{o}Q\frac{E[R_{t}]}{t}+C_{h}x-\frac{C_{h}}{2}mt+C_{h}Q\cdot\frac{1}{t}\int_{0}^{t}E[R_{s}]\,ds.$
We have
$C_{o}Q\frac{E[R_{t}]}{t}\to C_{o}m,$
and
$C_{h}Q\cdot\frac{1}{t}\int_{0}^{t}E[R_{s}]\,ds=C_{h}Q\cdot t\cdot\frac{1}{t^{2}}\int_{0}^{t}E[R_{s}]\,ds\to C_{h}Q\cdot t\cdot\frac{m}{2Q}=\frac{C_{h}}{2}mt.$
Thus, the terms proportional to $t$ cancel, and we obtain
$\lim_{t\to\infty}\frac{1}{t}E[C_{\mathrm{total}}(a,Q,t)]=C_{o}m+C_{h}x=C_{o}\bigl(\mu+\alpha\lambda+\lambda^{\prime}E[J_{1}]\bigr)+C_{h}x.$
Since the limit does not contain $Q$, the long-run average cost does not depend on $Q$ (nor on $a$). ∎
Corollary 3.18.
If $J_{k}$ follows an exponential distribution with parameter $\eta$,
(31) $\lim_{t\to\infty}\frac{1}{t}E[C_{\mathrm{total}}(a,Q,t)]=C_{o}\left(\mu+\alpha\lambda+\frac{\lambda^{\prime}}{\eta}\right)+C_{h}x.$
If $J_{k}$ follows a Gamma distribution with parameters $\beta,\eta$,
(32) $\lim_{t\to\infty}\frac{1}{t}E[C_{\mathrm{total}}(a,Q,t)]=C_{o}\left(\mu+\alpha\lambda+\frac{\beta\lambda^{\prime}}{\eta}\right)+C_{h}x.$
Acknowledgements
The author is deeply grateful to Professor Seiji Hiraba and Takumu Ooi for their continuous support, insightful suggestions, and invaluable guidance. The author also thanks Professor Aya Ishigaki and Yuriko Ono for fruitful collaboration and stimulating discussions.